how to put : 5x^2+9y^2+1 into standard form for ellipse
Answer by lwsshak3(11628)

how to put : 5x^2+9y^2+1 into standard form for ellipse
..
standard form for ellipse, x^2/a^2+y^2/b^2=1
given equation :5x^2+9y^2+1
=x^2/(1/5)+y^2/(1/9)=1
a^2=1/5
a=sqrt(1/5)
b^2=(1/9)
b=(1/3)
Since the denominator,(1/5), of the x^2 term is greater than the denominator,(1/9), of the y^2 term, this is an ellipse with a horizontal major axis.Its center is at (0,0).